GIVEN THAT LOGa(x)=3.58 AND LOGa(y)= 4.79, FIND LOGa(y/x).
Answer by joyofmath(189)

GIVEN THAT LOGa(x)=3.58 AND LOGa(y)= 4.79, FIND LOGa(y/x).


LOGa(y/x) = LOGa(y) - LOGa(x) = 4.79-3.58 = 1.21.